$SPAD/src/input fname.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool fname.output
)set message test on
)set message auto off
)clear all
)system touch /tmp/reado
)system chmod +r,-w /tmp/reado
)system touch /tmp/writo
)system chmod +w,-r /tmp/writo
--S 1 of 9
nullo: FNAME := new("", "nullo", "x")
--R
--R
--I   (1)  "nullo1406.x"
--R                                                               Type: FileName
--E 1
--S 2 of 9
reado: FNAME := filename("/tmp", "reado", "")
--R
--R
--R   (2)  "/tmp/reado"
--R                                                               Type: FileName
--E 2
--S 3 of 9
writo: FNAME := "/tmp/writo"
--R
--R
--R   (3)  "/tmp/writo"
--R                                                               Type: FileName
--E 3
--S 4 of 9
[directory reado, name reado, extension reado]
--R
--R
--R   (4)  ["/tmp","reado",""]

--R                                                           Type: List(String)
--E 4
--S 5 of 9
[directory writo, name writo, extension writo]
--R
--R
--R   (5)  ["/tmp","writo",""]
--R                                                           Type: List(String)
--E 5
--S 6 of 9
[directory nullo, name nullo, extension nullo]
--R
--R
--I   (6)  ["","nullo1406","x"]
--R                                                           Type: List(String)
--E 6
--S 7 of 9
[exists? reado, readable? reado, writable? reado]
--R
--R
--R   (7)  [true,true,false]
--R                                                          Type: List(Boolean)
--E 7
--S 8 of 9
[exists? writo, readable? writo, writable? writo]
--R
--R
--R   (8)  [true,true,true]
--R                                                          Type: List(Boolean)
--E 8
--S 9 of 9
[exists? nullo, readable? nullo, writable? nullo]
--R
--R
--R   (9)  [false,false,true]
--R                                                          Type: List(Boolean)
--E 9
)system rm -f /tmp/reado
)system rm -f /tmp/writo
)spool
)lisp (bye)

———-

References^{[1] nothing}

5